FAR-INFRARED LASER MAGNETIC RESONANCE SPECTROSCOPIC STUDY OF THE ν_{2} BENDING FUNDAMENTAL OF THE CCN RADICAL IN ITS X̃^{2}Π_{r} STATE

\underline{M. D. ALLEN}, K. M. EVENSON, Time and Frequency Division, National Institute of Standards and Technology, Boulder, CO 80303-3328; and J. M. BROWN, The Physical Chemistry Laboratory, Oxford University, South Parks Road, Oxford OX1 3QZ, United Kingdom.

Vibration-rotation transitions between the (010) μ^{2}Σ^{−} - (000) X̃^{2}Π_{r} vibronic states, were recorded using far-infrared laser magnetic resonance (FIR-LMR) spectroscopic techniques. These transitions occur near 200 cm^{−1}for the (010) μ^{2}Σ^{−} - (000) X̃^{2}Π_{1/2} transition and 160 cm^{−1}for the (010) μ^{2}Σ^{−} - (000) X̃^{2}Π_{3/2} transition. This is the first direct measurement of the ν_{2} band of CCN and with the inclusion of optical data, similar to the one found in the paper by Kohguchi et al.,^{a}, microwave data from Ohshima and Endo^{b}, and spin-orbit data^{c} a least squares fit has resulted in an accurate determination of molecular parameters for the ground and the (010) μ^{2}Σ^{−} states. The data were fit using an N^{2} effective Hamiltonian modified to include the Renner-Teller effect.

aH. Kohguchi, Y. Ohshima, Y. Endo, J . Chem. Phys. 106, 5429 (1997).

bY. Ohshima and Y. Endo, J .Molec.Spectros. 172, 225-232 (1995).

cCalculated from data reported in M. Kakimoto and T. Kasuya, J .Molec.Spectros. 94, 380-392 (1982) and

K. Kawaguchi, T. Suzuki, S. Saito, and E. Hirota, J .Molec.Spectros. 106, 320-329 (1984).